# RoboSpatial: Teaching Spatial Understanding to 2D and 3D Vision-Language Models for Robotics

### v1

@@ 5 Conclusion
We introduce RoboSpatial and RoboSpatial-Home, a large-scale 2D/3D spatial understanding training and evaluation dataset tailored for robotics. Experimental results show that models trained with RoboSpatial are able to understand spatial relationships, generalize to unseen relationships, and infer nuanced reference frames, making them applicable in a wide range of tasks that require spatial understanding. We further demonstrate the real-world applicability of RoboSpatial with robot experiments. Furthermore, our automatic data generation pipeline can be used to extend the dataset whenever needed. We show that RoboSpatial has the potential to serve as a foundation for broader applications in robotic scenarios that require spatial understanding.
Acknowledgement
The authors would like to thank Youngsun Wi, Hyunho Ahn, Minjae Bae, and Hojin Yoo for providing images in the RoboSpatial-Home dataset. Our research is supported in part by ARL W911NF2220144 and resources from the Ohio Supercomputer Center. These research findings were partially derived using the Matterport dataset, the use of which is governed by its terms of service: https://kaldir.vc.in.tum.de/matterport/MP_TOS.pdf.
@@ Acknowledgement
Appendix A Dataset Details
A.1 Dataset Statistics
@@
We provide the full dataset statistics in Tab. 8. For all training, we use only 900,000 spatial relationships, equally divided across all datasets, due to computational constraints. We further experiment on the effect of data scaling on Tab. 9 and explain the results. Notably, HOPE [50] and GraspNet-1B [15] contain similar tabletop images captured from different perspectives, resulting in lower dataset diversity for the tabletop environment. We plan to enhance the diversity of our dataset by incorporating additional tabletop datasets.
A.2 Dataset Generation Details
Frame Generation. We explain how answers are generated for each frame. Each question type has three answers from ego-centric, object-centric, and world-centric perspectives. Each frame may share the same answer since not all frames lead to unique answers. For our answer generation, we used the following logic:
  • Ego-centric: The default perspective from the ego-centric view.
  • Object-centric: Using the oriented bounding box directions, we determine the front of the object. With this information, we assign front, behind, left, and right. Above and below remain the same as in the ego-centric perspective.
  • World-centric: Using the z-coordinate of the oriented bounding box, we modify the above and below relationships to reflect whether an object is above or below another object with respect to elevation.
Compatibility Generation. For compatibility, we construct a top-down map as shown in Fig. 5. Using the top-down map and the top-down 2D bounding box of the object to be placed, we determine if there exists an empty space within a threshold distance (i.e., 1% of the object’s longest width or length). If the object meets this condition, it is deemed compatible to be fitted.
@@ A.2 Dataset Generation Details
Validation 5 scans 100 100 100 100
Test 5 scans 100 100 100 100
ScanNet [11] Train 1512 scans 278402 298439 298439 298439
Validation 12 scans 1200 500 500 500
Test 12 scans 1200 500 500 500
3RScan [52] Train 1543 scans 365355 298439 298439 298439
Validation 9 scans 900 400 400 400
Test 9 scans 900 400 400 400
Tabletop HOPE [50] Train 50 scenes 50000 36317 36317 36317
Validation 10 scenes 50 500 500 500
Test 47 scenes 235 500 500 500
GraspNet-1B [15] Train 100 scenes 25500 36317 36317 36317
@@
Appendix B Implementation Details
B.1 Model Training
VILA [30] We initialize our model from Efficient-Large-Model/Llama-3-VILA1.5-8B on Hugging Face. We use the fine-tuning script from the VILA GitHub repository to train our model using the default hyperparameters.
LLaVA-NeXT [34] We initialize our model from lmms-lab/llama3-llava-next-8b on Hugging Face. We use the LLaVA-Next fine-tuning script from the LLaVA-Next repository using the default hyperparameters.
SpaceLLaVA [8] As official code and weights for SpatialVLM [8] is not released, we use a community implementation which is endorsed by SpatialVLM [8] authors. We initialize our model from remyxai/SpaceLLaVA from Hugging Face. We use LLaVA-1.5 finetuning script from LLaVa [35] repository using the default hyperparameters.
RoboPoint [57] We initialize our model from wentao-yuan/robopoint-v1-vicuna-v1.5-13b on Hugging Face. We use the fine-tuning script provided in the RoboPoint [57] GitHub repository to train our model using the default hyperparameters.
3D-LLM [18] We initialize our model using the pretrain_blip2_sam_flant5xl_v2.pth checkpoint downloaded from the official GitHub repository. Since the model requires preprocessing of multiview images, we follow the author’s pipeline to process multiview images from our environments. Because the model does not accept image input, we append the following text in front of the question to ensure the model understands the perspective from which the question is being asked: “I am facing anchor object.” We use the default hyperparameters and train the model for 20 epochs per the author’s guidelines. We choose the best model based on validation accuracy.
LEO [20] We initialize our model from the sft_noact.pth checkpoint downloaded from the official GitHub repository.
Annotation Size 100K 300K 900k (Default) 1.8M 3M (Full)
LLaVa-Next [34] 38.1 46.7 60.5 65.8 72.4
Table 9: Results of scaling experiment on LLaVa-Next [34] with varied spatial relationship annotations. Average accuracy on held-out test set is reported.
B.2 Robot Setup
For picking, we find which object the point maps to using SAM 2 [42] and execute our picking behavior on that object. For placing, we simply compute the 3D coordinate based on the depth value at that pixel and place the object at that coordinate. There were no failures due to cuRobo [47] failing. The experiments were purposely designed to consist of behaviors that our robot system can handle in order to avoid introducing irrelevant factors. The picking behavior consists of computing a top-down grasp pose and reaching it with cuRobo [47]. To compute the grasp pose:
  • 1.
  • 2.
    The grasp orientation is orthogonal to the major axis.
  • 3.
@@ B.2 Robot Setup
Table 10: Results of per frame accuracy of existing 2D/3D VLMs on a held-out test split of images and scans. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset. The best result for each column is bolded.
Appendix C Full Results
C.1 Data Scaling
In Tab. 9, we experiment with scaling the number of annotations while keeping images fixed. We found that even though the number of images stays consistent, increasing the number of annotations can improve performance. For future work, we plan to apply our data generation pipeline to a diverse set of indoor and tabletop environments to further improve the performance of our models.
C.2 Accuracy Per Frame
@@
We show the results per frame in Tab. 10 for our out-of-domain test set. From the results, we can see a distinct difference between 2D and 3D VLMs in understanding the world-centric frame before training with RoboSpatial. Baseline 2D VLMs have trouble understanding the world-centric frame, which involves understanding elevation, while 3D VLMs comparatively excel at it. Furthermore, we can see that since baseline 3D VLMs are trained on point clouds without information of perspective, their accuracy in ego-centric and object-centric frames is lower. However, with RoboSpatial training, we were able to teach the 3D VLMs to think in a certain frame, thus considerably improving their performance on ego-centric and object-centric frames. However, we hypothesize that, due to their design—specifically, the lack of a means to visually inject perspective information since they require complete 3D point clouds—3D VLMs still lag behind 2D VLMs on ego-centric and object-centric frames.
C.3 Robot Experiments
@@
We present additional results from our robot experiments in Fig. 6. We observe that models trained with RoboSpatial consistently outperform baseline models in most cases, even though the prompt is not optimized for RoboSpatial-trained models. This demonstrates that the power of VLMs enables templated language to generalize to language unseen during training while maintaining spatial understanding capabilities. However, even with RoboSpatial training, the models struggle with understanding stacked items, indicating a need for further data augmentation with diverse layouts. In a few cases, RoboSpatial training adversely affects performance, especially with RoboPoint [57]. We hypothesize that mixing the dataset with RoboPoint training data and RoboSpatial training data may lead to unforeseen side effects, particularly in grounding objects. Nevertheless, we demonstrate that RoboSpatial training enhances VLM’s spatial understanding in real-life robotics experiments, even with freeform language.
@@
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57]
@@
Molmo [12]
LLaVa-Next [34]
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57]
@@
S-LLaVa-Next [34]
@@
S-RoboPoint [57]
Question: is there an object that is not in a stack?
@@
LLaVa-Next [34]
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Question: is there space to place one of the cans on the cheez-it box?
@@
S-RoboPoint [57] ×
Question: pick the highest object on the stack of two objects
@@
S-RoboPoint [57] ×
Figure 6: Additional robot experiments. A green check mark indicates that the model answered correctly. The S- prefix denotes a model trained with RoboSpatial. The questions are purposely not cleaned to reflect realistic language inputs.
Appendix D More Qualitative Examples
Fig. 7 and Fig. 8 present additional qualitative comparisons between models trained on RoboSpatial. Our findings demonstrate that models trained on RoboSpatial consistently exhibit spatial understanding in the challenging RoboSpatial-Home dataset, even outperforming closed models like GPT-4o [40]. However, we observed that object grounding is a crucial prerequisite for spatial understanding; the improvement is often hindered by the model’s inability to ground objects in cluttered scenes, where GPT-4o performs more effectively. Additionally, in Fig. 8, we show that the RoboSpatial-trained model successfully generalizes to unseen spatial relationships in Blink-Spatial [17], including those involving distance, such as ”touching.”
Figure 7: Qualitative results on RoboSpatial-Home. For spatial context questions, we omit the baseline models which fail to generate a point inside the image. Furthermore, the first predict point for each model is shown in the image.
Figure 8: Qualitative results on Blink-Spatial [17]. RoboSpatial-trained model can generalize to unseen spatial relationships.

### v2

@@ 5 Conclusion
We introduce RoboSpatial, RoboSpatial-Val, and RoboSpatial-Home, a large-scale 2D/3D spatial understanding training and evaluation dataset tailored for robotics. Experimental results show that models trained with RoboSpatial are able to understand spatial relationships, generalize to unseen relationships, and infer nuanced reference frames, making them applicable in a wide range of tasks that require spatial understanding. We further demonstrate the real-world applicability of RoboSpatial with robot experiments. In addition, our automatic data generation pipeline can be used to extend the dataset to new data sources and spatial relations. We show that RoboSpatial has the potential to serve as a foundation for broader applications in robotics which require spatial understanding.
Acknowledgements
The authors thank Youngsun Wi, Hyunho Ahn, Hojin Yoo, and Minjae Bae for providing images in the RoboSpatial-Home dataset. Our research was supported in part by ARL W911NF2220144 and resources from the Ohio Supercomputer Center. These research findings were partially derived using the Matterport dataset, the use of which is governed by: https://kaldir.vc.in.tum.de/matterport/MP_TOS.pdf.
@@ Acknowledgements
Appendix A Limitations
While RoboSpatial significantly improves spatial reasoning capabilities in VLMs, certain design choices naturally introduce trade-offs and areas for future exploration.
First, the dataset relies on a top-down occupancy map to identify and annotate empty regions for spatial context and compatibility tasks. This approach simplifies reasoning about object placement on horizontal surfaces and enables efficient data generation, but it currently does not support spatial questions involving containment—such as whether an object can fit inside or under another object—which would require more detailed volumetric modeling.
Second, although the models are deployed on a real robot using a modular approach, we do not yet explore tighter forms of integration such as training it jointly with robot trajectories [27]. Investigating these alternatives could enhance downstream policy learning and enable more seamless end-to-end systems.
Finally, RoboSpatial focuses on indoor and tabletop scenes containing objects commonly encountered in household environments, and does not include humans or animals. This reflects the nature of source datasets and our emphasis on robot object manipulation. While this limits coverage of social or dynamic interaction scenarios, trained models still generalizes well to out-of-distribution benchmarks like BLINK, which include humans and animals—suggesting that the learned spatial representations are broadly transferable.
Appendix B Dataset Details
B.1 Dataset Statistics
@@
We provide the full dataset statistics in Tab. 7. For all training, we use only 900,000 spatial relationships, sampled equally across all datasets, due to computational constraints. We further experiment on the effect of data scaling on Tab. 9 and explain the results. Notably, HOPE [54] and GraspNet-1B [12] contain similar tabletop images captured from different perspectives, resulting in lower dataset diversity for the tabletop environment. We plan to enhance the diversity of RoboSpatial by incorporating additional tabletop datasets.
B.2 Choice of Spatial Relationships
In designing the dataset, we focused on spatial relationships that directly impact robotic perception, planning, and interaction: context, compatibility, and configuration. These were selected to reflect the core spatial reasoning challenges that robots encounter when operating in complex, real-world environments.
We intentionally excluded tasks such as object counting, as we consider them to fall outside the scope of spatial understanding. While counting is an important visual reasoning skill, it does not require reasoning about spatial relations between objects or between objects and their environment. For example, determining that “three cups are on the table” is a perceptual task rather than a spatial reasoning one. As such, counting may complement but does not substitute for the types of relational reasoning we target. We leave the integration of counting tasks into spatial benchmarks as future work.
Similarly, we exclude tasks that rely solely on distance measurements. Although distance is a fundamental spatial quantity, it is difficult to define consistently across different environments, object scales, and robot embodiments. Absolute distances can vary significantly between indoor and outdoor scenes, small and large objects, or different robot perspectives, making them hard to normalize or interpret in a general way. Moreover, distance alone often lacks the relational semantics required for higher-level reasoning—for example, understanding that an object is behind, above, or in front of others. RoboSpatial instead focuses on spatial relationships that are more invariant, interpretable, and transferable across diverse robotic scenarios.
That said, the data generation pipeline is general and could readily support auxiliary tasks involving object counting or distance estimation if desired. These metrics may serve as useful complements in future extensions of the benchmark or as auxiliary supervision signals in model training.
B.3 Object Grounding Dataset
To support accurate spatial understanding, we generate an auxiliary dataset for object grounding. Many spatial reasoning tasks assume that the model can correctly identify which object is being referred to in the scene. However, in practice, this can be a major source of error—especially in cluttered environments or when multiple instances of the similar object type are present.
The grounding dataset provides direct supervision to help models learn to associate text descriptions with specific objects in the image. For each image, we include a set of object descriptions (e.g., “the keyboard” or “the chair”) paired with the corresponding 2D bounding box of the object in the image. These 2D boxes are projected from the annotated 3D bounding boxes using camera intrinsics and extrinsics.
A total of 100k grounding QA pairs are generated and used during training to reduce reference ambiguity and improve object identification accuracy in spatial tasks. While not part of the main spatial reasoning taxonomy, grounding accuracy is a prerequisite for answering spatial questions correctly, and we find that including this data helps reduce errors caused by incorrect object identification.
@@ B.3 Object Grounding Dataset
Validation 10 scans 200 200 200 200
ScanNet [8] Train 1514 scans 280402 299039 299039 299039
Validation 12 scans 400 400 400 400
3RScan [56] Train 1543 scans 366755 298839 298839 298839
Validation 18 scans 400 400 400 400
Tabletop HOPE [54] Train 60 scenes 50050 36817 36817 36817
Validation 47 scenes 235 500 500 500
GraspNet-1B [12] Train 130 scenes 25620 36817 36817 36817
@@
B.4 Dataset Generation Details
The dataset generation pipeline is detailed in the main text (subsection 3.2), which introduces a two-stage process for computing 3D spatial relationships and projecting them into 2D image space. Here, we expand on implementation details not covered in the main paper and provide clarification on the reasoning logic used in spatial annotation.
Reference Frame Annotation. For each spatial configuration question, we label relationships from three perspectives: ego-centric (camera view), object-centric (based on object heading), and world-centric (aligned with the dataset’s global frame). To compute object-centric directions, we use the heading vector of each oriented 3D bounding box to define the “front” of the object. Left, right, behind, and front relations are then assigned accordingly. World-centric annotations modify vertical relationships (above/below) using global $z$-coordinates to reflect elevation.
Surface Detection and Free Space Sampling. To identify support surfaces such as tables, counters, or floors, we use GPT-4o to select candidate objects that are likely to support placement. A top-down occupancy map is constructed from bounding boxes in the scene  Fig. 5. We sample 3D points in unoccupied regions and project them into the image plane for spatial context tasks. Points are filtered via occlusion checks using raycasting, ensuring sampled points are visible and unobstructed.
Compatibility Check and Object Placement. For spatial compatibility, we simulate placing a virtual object bounding box at candidate locations. The placement must fit without intersecting other objects and must allow a clearance of at least 10 cm in all axes. We allow in-plane rotation and translation to test flexible placement. This provides a binary label (True/False) indicating whether the object can be compatibly placed in the region.
Output Format. Though RoboSpatial uses point prediction for ease of integration with robot setups, the pipeline also supports mask-based outputs and can be extended in future work.
Appendix C Implementation Details
@@
Table 8: Results of per frame accuracy of existing 2D/3D VLMs on a RoboSpatial-Val. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on RoboSpatial. The best result for each column is bolded.
Annotation Size 100K 300K 900k (Default) 1.8M 3M (Full)
LLaVa-Next [35] 38.1 46.7 60.5 65.8 72.4
Table 9: Results of scaling experiment on LLaVa-Next [35] with varied number of spatial relationship annotations. Average accuracy on RoboSpatial-Val is reported.
MMMUval MMEp MMEc MMBenchdev
LLaVA-NeXT 39.4 1561.8 305.4 71.6
+RoboSpatial 39.8 1604.5 293.2 71.6
Table 10: Evaluation on general-purpose multimodal benchmarks (MMMU, MME, MMBench) to assess whether training on RoboSpatial affects commonsense and factual reasoning.
Base Auxiliary RoboSpatial Both
LLaVA-NeXT 30.3 32.4 51.8 60.5
Table 11: Ablation study evaluating the impact of the auxiliary grounding dataset on RoboSpatial-Val.
C.1 Model Training
C.2 Model Setup
VILA [30] We initialize the model from Efficient-Large-Model/Llama-3-VILA1.5-8B on Hugging Face. We use the fine-tuning script from the VILA GitHub repository to train the model using the default hyperparameters.
LLaVA-NeXT [35] We initialize the model from lmms-lab/llama3-llava-next-8b on Hugging Face. We use the LLaVA-Next fine-tuning script from the LLaVA-Next repository using the default hyperparameters.
SpaceLLaVA [5] As official code and weights for SpatialVLM [5] is not released, we use a community implementation which is endorsed by SpatialVLM [5] authors. We initialize the model from remyxai/SpaceLLaVA from Hugging Face. We use LLaVA-1.5 finetuning script from LLaVa [34] repository using the default hyperparameters.
RoboPoint [62] We initialize the model from wentao-yuan/robopoint-v1-vicuna-v1.5-13b on Hugging Face. We use the fine-tuning script provided in the RoboPoint [62] GitHub repository to train the model using the default hyperparameters.
3D-LLM [18] We initialize the model using the pretrain_blip2_sam_flant5xl_v2.pth checkpoint downloaded from the official GitHub repository. Since the model requires preprocessing of multiview images, we follow the author’s pipeline to process multiview images from the environments. Because the model does not accept image input, we append the following text in front of the question to ensure the model understands the perspective from which the question is being asked: “I am facing anchor object.” We use the default hyperparameters and train the model for 20 epochs per the author’s guidelines. We choose the best model based on validation accuracy.
LEO [20] We initialize the model from the sft_noact.pth checkpoint downloaded from the official GitHub repository.
Appendix D More Results
D.1 Accuracy Per Reference Frame
@@
We show the results per frame in Tab. 8 for RoboSpatial-Val. From the results, we can see a distinct difference between 2D and 3D VLMs in understanding the world-centric frame before training with RoboSpatial. Baseline 2D VLMs have trouble understanding the world-centric frame, which involves understanding elevation, while 3D VLMs comparatively excel at it. Furthermore, we can see that since baseline 3D VLMs are trained on point clouds without information of perspective, their accuracy in ego-centric and object-centric frames is lower. However, with RoboSpatial training, we were able to teach the 3D VLMs to think in a certain frame, thus considerably improving their performance on ego-centric and object-centric frames. However, we hypothesize that, due to their design—specifically, the lack of a means to visually inject perspective information since they require complete 3D point clouds—3D VLMs still lag behind 2D VLMs on ego-centric and object-centric frames.
D.2 Data Scaling
In Tab. 9, we experiment with scaling the number of annotations while keeping images fixed. We found that even though the number of images stays consistent, increasing the number of annotations can improve performance. For future work, we plan to apply the data generation pipeline to a diverse set of indoor and tabletop environments to further improve the performance of the models.
D.3 Commonsense Knowledge Retention
To ensure that training on RoboSpatial does not degrade a model’s general reasoning or commonsense capabilities, we evaluate the RoboSpatial-trained model on a suite of standard multimodal benchmarks: MMMU [63], MME [14], and MMBench [36]. As shown in Table 10, the RoboSpatial-trained model maintains or slightly improves performance across all benchmarks, suggesting that spatial fine-tuning preserves broader knowledge capabilities.
D.4 Ablation of the Auxiliary Grounding Dataset
As shown in Table 11, training on the auxiliary dataset alone yields a small improvement over the base model (+2.1), but it falls far short of the gains achieved with RoboSpatial, which is explicitly designed to teach spatial reasoning. This confirms that grounding supervision alone is insufficient for spatial understanding. However, combining both datasets leads to the best performance, suggesting that improving object localization can complement spatial supervision when jointly trained.
D.5 Robot Experiments Details
D.5.1 Robot Setup
For picking, we find which object the point maps to using SAM 2 [45] and execute the picking behavior on that object. For placing, we simply compute the 3D coordinate based on the depth value at that pixel and place the object at that coordinate. There were no failures due to cuRobo [52] failing. The experiments were purposely designed to consist of behaviors that our robot system can handle in order to avoid introducing irrelevant factors. The picking behavior consists of computing a top-down grasp pose and reaching it with cuRobo [52]. To compute the grasp pose:
  1. 1.
  2. 2.
    The grasp orientation is orthogonal to the major axis.
  3. 3.
D.5.2 Additional Results
@@
We present additional results from the robot experiments in Fig. 6. We observe that models trained with RoboSpatial consistently outperform baseline models in most cases, even though the prompt is not optimized for RoboSpatial-trained models. This demonstrates that the power of VLMs enables templated language to generalize to language unseen during training while maintaining spatial understanding capabilities. However, even with RoboSpatial training, the models struggle with understanding stacked items, indicating a need for further data augmentation with diverse layouts. In a few cases, RoboSpatial training adversely affects performance, especially with RoboPoint [62]. We hypothesize that mixing the dataset with RoboPoint training data and RoboSpatial training data may lead to unforeseen side effects, particularly in grounding objects. Nevertheless, we demonstrate that RoboSpatial training enhances VLM’s spatial understanding in real-life robotics experiments, even with freeform language.
@@
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
Molmo [9]
@@
LLaVa-Next [35]
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
@@
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62]
RoboPoint-FT [62]
Molmo [9]
@@
Question: is there an object that is not in a stack?
LLaVa-Next [35]
LLaVa-Next-FT [35]
RoboPoint [62]
RoboPoint-FT [62]
@@
GPT-4o [42]
Question: is there space to place one of the cans on the cheez-it box?
@@
LLaVa-Next-FT [35] ×
RoboPoint [62] ×
RoboPoint-FT [62] ×
Molmo [9] ×
GPT-4o [42] ×
Question: pick the highest object on the stack of two objects
LLaVa-Next [35] ×
LLaVa-Next-FT [35] ×
RoboPoint [62] ×
RoboPoint-FT [62] ×
Molmo [9] ×
GPT-4o [42] ×
Figure 6: Additional robot experiments. A green check mark indicates that the model answered correctly. The -FT suffix denotes a model trained with RoboSpatial. The questions are purposely not cleaned to reflect realistic language inputs.
D.6 More Qualitative Examples
Fig. 7 present additional qualitative comparisons between models trained on RoboSpatial. The findings demonstrate that models trained on RoboSpatial consistently exhibit spatial understanding in the challenging RoboSpatial-Home dataset, even outperforming closed models like GPT-4o [42]. However, we observed that object grounding is a crucial prerequisite for spatial understanding; the improvement is often hindered by the model’s inability to ground objects in cluttered scenes, where GPT-4o performs more effectively. Additionally, we show that the RoboSpatial-trained model successfully generalizes to unseen spatial relationships in BLINK-Spatial [15], including those involving distance, such as ”touching.”
Question: Pinpoint several points within the vacant space situated to the left of the pot.
Answer
LLaVa-Next [35]
LLaVa-Next-FT [35]
@@ D.6 More Qualitative Examples
RoboPoint-FT [62]
Molmo [9]
@@
Question: Pinpoint several points within the vacant space situated behind the trash bin.
Answer
LLaVa-Next [35]
LLaVa-Next-FT [35]
RoboPoint [62]
RoboPoint-FT [62]
@@
GPT-4o [42]
Question: Can the lamp fit in front of the shelf?
Answer Yes
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
Molmo [9] ×
@@
Question: Is the lamp above the shelf?
Answer Yes
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
@@
GPT-4o [42]
Question: Is the dining table touching the donut?
Answer Yes
@@
RoboPoint-FT [62]
Molmo [9] ×
@@
Question: Can the pot fit above the fridge?
Answer Yes
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
@@
RoboPoint-FT [62]
Molmo [9] ×
GPT-4o [42] ×
Question: Is the couch under the suitcase?
Answer Yes
LLaVa-Next [35] ×
LLaVa-Next-FT [35]
RoboPoint [62] ×
RoboPoint-FT [62]
Molmo [9] ×
GPT-4o [42] ×
Figure 7: Qualitative results on spatial reasoning benchmarks. The -FT suffix denotes a model trained with RoboSpatial. The first three rows show examples from RoboSpatial-Home, covering spatial context, spatial compatibility, and spatial configuration. For spatial context questions, only the first predicted point from each model is shown. The fourth row shows generalization to unseen spatial relationships on the Blink-Spatial [15] dataset, demonstrating that the RoboSpatial-trained model can transfer to unseen relationships.